# Spectrum of Schrödinger operators on subcovering graphs

### v2

@@ 1. Introduction
This paper was partly motivated by the nice article [KP07] describing simple connections between spectra of Schrödinger operators on graphene and nanotubes. Graphene is a single 2D layer of graphite forming a hexagonal lattice. The hexagonal lattice is invariant under translations by vectors $\mathfrak{a}_{1}$ and $\mathfrak{a}_{2}$, see Fig. 1a. A carbon nanotube is a hexagonal lattice ”rolled up” into a cylinder. In carbon nanotubes, the graphene sheet is ”rolled up” in such a way that a so-called chiral vector $\mathfrak{t}=t_{1}\mathfrak{a}_{1}+t_{2}\mathfrak{a}_{2}$, where $t:=(t_{1},t_{2})\in{\mathbb{Z}}^{2}$, becomes the circumference of the tube, see Fig. 1b,c. In [KP07], applying a simple restriction procedure to the dispersion relation for the Schrödinger operator on the graphene lattice, the authors described the spectra of the Schrödinger operators on any carbon nanotube: zigzag ($t=(n,0)$), armchair ($t=(n,n)$) or chiral. In particular, they showed that the spectra of the Schrödinger operators on a nanotube and graphene coincide as sets if and only if $t_{1}-t_{2}$ is divisible by 3, and these spectra are asymptotically close as the length of the chiral vector $\mathfrak{t}$ tends to infinity. The similar approach was used in [D15] to derive spectral properties of graphyne nanotubes from the dispersion relation of the graphyne lattice.
@@
Let ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ be a connected infinite graph, possibly having loops and multiple edges and embedded into ${\mathbb{R}}^{d}$. Here ${\mathcal{V}}$ is the set of its vertices and ${\mathcal{E}}$ is the set of its (unoriented) edges. An edge from ${\mathcal{E}}$ between vertices $u,v\in{\mathcal{V}}$ will be denoted as the (unordered) pair $\{u,v\}$ and said to be incident to the vertices $u$ and $v$.
@@
Remark 1.1.
A periodic graph ${\mathcal{G}}$ could be defined as an abstract infinite graph equipped with an action of a finitely generated free abelian group (a lattice) $\Gamma$ (see, e.g., Chapter 4 in [BK13]). The embedding of ${\mathcal{G}}$ into ${\mathbb{R}}^{d}$ is just a simple geometric model (realization) of periodic graphs.
@@
We consider the Schrödinger operator $H$ acting on $\ell^{2}({\mathcal{V}})$ and given by
$H=\Delta+Q,$ (1.1)
where $\Delta$ is the discrete Laplacian having the form
$(\Delta f)(v)=\sum_{\{v,u\}\in{\mathcal{E}}}\big{(}f(v)-f(u)\big{)},\qquad f\in\ell^{2}({\mathcal{V}}),\qquad v\in{\mathcal{V}},$ (1.2)
@@
The sum in (1.2) is taken over all edges $\{v,u\}\in{\mathcal{E}}$ incident to the vertex $v$. It is known that the Schrödinger operator $H$ is a bounded self-adjoint operator on $\ell^{2}({\mathcal{V}})$ (see, e.g., [SS92]).
@@
$f(v+t_{1}\mathfrak{a}_{1}+\ldots+t_{d}\mathfrak{a}_{d})=e^{i\langle t,k\rangle}f(v),\qquad\forall\,v\in{\mathcal{V}},\qquad\forall\,t=(t_{1},\ldots,t_{d})\in{\mathbb{Z}}^{d},$ (1.3)
@@
which, in particularly, yields
@@
Some of $\lambda_{j}(\cdot)$ may be constant, i.e., $\lambda_{j}(\cdot)=\Lambda=\mathop{\mathrm{const}}\nolimits$, on some subset of ${\mathcal{B}}$ of positive Lebesgue measure. In this case the Schrödinger operator $H$ on ${\mathcal{G}}$ has the eigenvalue $\Lambda$ of infinite multiplicity. We call $\Lambda$ a flat band. Thus, the spectrum of the Schrödinger operator $H$ on a periodic graph ${\mathcal{G}}$ has the form
@@
i) The first spectral band $\sigma_{1}(H)=[\lambda_{1}^{-},\lambda_{1}^{+}]$ is never a flat band and $\lambda_{1}^{-}=\lambda_{1}(0)$, see, e.g., Theorem 2.1.ii in [KS19].
@@ 1.2. Subcovering graphs
Let $\Gamma$ be a lattice of rank $d\geqslant 2$ in ${\mathbb{R}}^{d}$ with a basis $\{\mathfrak{a}_{1},\ldots,\mathfrak{a}_{d}\}$, and let $\Gamma_{\mathfrak{t}}$ be a sublattice of $\Gamma$ with a basis $\mathfrak{t}:=\{\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}\}\subset\Gamma$, where $d_{o}<d$.
We consider a subcovering graph ${\mathcal{G}}_{\mathfrak{t}}$ of a $\Gamma$-periodic graph ${\mathcal{G}}$ obtained as the quotient of ${\mathcal{G}}$ with respect to the sublattice $\Gamma_{\mathfrak{t}}$, i.e.,
@@
The hexagonal lattice ${\bf G}$ is a $\Gamma$-periodic graph, where $\Gamma\subset{\mathbb{R}}^{2}$ is the lattice with the basis $\mathfrak{a}_{1},\mathfrak{a}_{2}$, see Fig. 1a. For non-zero vectors $\mathfrak{t}\in\Gamma$ the subcoverings ${\bf G}_{\mathfrak{t}}={\bf G}/{\mathbb{Z}}\mathfrak{t}$ of the hexagonal lattice ${\bf G}$ are known as nanotubes. For example, the nanotube ${\bf G}_{\mathfrak{t}}$ with the chiral vector $\mathfrak{t}=2\mathfrak{a}_{1}+\mathfrak{a}_{2}$ is shown in Fig. 1b. The so-called zigzag nanotube ${\bf G}_{\mathfrak{t}}$, $\mathfrak{t}=2\mathfrak{a}_{1}$, is presented in Fig. 1c.
@@
Figure 2. a) The cubic lattice ${\mathbb{L}}^{3}$; the blue vectors $\mathfrak{a}_{1},\mathfrak{a}_{2},\mathfrak{a}_{3}$ are the periods of ${\mathbb{L}}^{3}$; the fundamental cell $\Omega=[0,1)^{3}$ is shaded. The chiral vectors $\mathfrak{t}_{1}$ and $\mathfrak{t}_{2}$ are marked in red color. b) The subcovering ${\mathbb{L}}^{3}_{\mathfrak{t}_{1}}$ of ${\mathbb{L}}^{3}$ with the chiral vector $\mathfrak{t}_{1}=\mathfrak{a}_{1}+\mathfrak{a}_{2}-\mathfrak{a}_{3}$ (the triangular lattice). c) The subcovering ${\mathbb{L}}^{3}_{\mathfrak{t}}$ of ${\mathbb{L}}^{3}$ with $\mathfrak{t}=\{\mathfrak{t}_{1},\mathfrak{t}_{2}\}$, where the chiral vectors $\mathfrak{t}_{1}=\mathfrak{a}_{1}+\mathfrak{a}_{2}-\mathfrak{a}_{3}$ and $\mathfrak{t}_{2}=2\mathfrak{a}_{1}+\mathfrak{a}_{2}$. d) The fundamental graph ${\mathbb{L}}^{3}_{*}$ of the cubic lattice ${\mathbb{L}}^{3}$ and any its subcovering ${\mathbb{L}}^{3}_{\mathfrak{t}}$ with a primitive set $\mathfrak{t}$ of chiral vectors.
@@
The cubic lattice ${\mathbb{L}}^{3}$ is a ${\mathbb{Z}}^{3}$-periodic graph, see Fig. 2a. Let $\{\mathfrak{a}_{1},\mathfrak{a}_{2},\mathfrak{a}_{3}\}$ be the standard basis of ${\mathbb{Z}}^{3}$. The subcovering ${\mathbb{L}}^{3}_{\mathfrak{t}_{1}}={\mathbb{L}}^{3}/{\mathbb{Z}}\mathfrak{t}_{1}$ of the cubic lattice ${\mathbb{L}}^{3}$ with the chiral vector $\mathfrak{t}_{1}=\mathfrak{a}_{1}+\mathfrak{a}_{2}-\mathfrak{a}_{3}$ is the triangular lattice, Fig. 2b. Let ${\mathbb{Z}}^{3}_{\mathfrak{t}}$ be the sublattice of ${\mathbb{Z}}^{3}$ with the basis $\mathfrak{t}:=\{\mathfrak{t}_{1},\mathfrak{t}_{2}\}$, where $\mathfrak{t}_{1}=\mathfrak{a}_{1}+\mathfrak{a}_{2}-\mathfrak{a}_{3}$ and $\mathfrak{t}_{2}=2\mathfrak{a}_{1}+\mathfrak{a}_{2}$. Then the subcovering ${\mathbb{L}}^{3}_{\mathfrak{t}}={\mathbb{L}}^{3}/{\mathbb{Z}}^{3}_{\mathfrak{t}}$ of the cubic lattice ${\mathbb{L}}^{3}$ is the ${\mathbb{Z}}$-periodic graph shown in Fig. 2c. Note that this ${\mathbb{Z}}$-periodic graph can also be considered as the subcovering of the triangular lattice (which is a ${\mathbb{Z}}^{2}$-periodic graph) with the chiral vector $\mathfrak{t}_{2}$.
@@
c)$v_{2}$$v_{1}$${\mathcal{G}}_{\mathfrak{t}}$
@@
The diamond lattice ${\mathcal{G}}$ shown in Fig. 3a is a $\Gamma$-periodic graph, where $\Gamma\subset{\mathbb{R}}^{3}$ is the lattice with the basis $\mathfrak{a}_{1},\mathfrak{a}_{2},\mathfrak{a}_{3}$. The subcovering ${\mathcal{G}}_{\mathfrak{t}}={\mathcal{G}}/{\mathbb{Z}}\mathfrak{t}$ of the diamond lattice ${\mathcal{G}}$ with the chiral vector $\mathfrak{t}=\mathfrak{a}_{1}+\mathfrak{a}_{2}-\mathfrak{a}_{3}$ is shown in Fig. 3b. Note that this subcovering of ${\mathcal{G}}$ is isomorphic to the square lattice. Let $\Gamma_{\mathfrak{t}}$ be the sublattice of $\Gamma$ with the basis $\mathfrak{t}:=\{\mathfrak{t}_{1},\mathfrak{t}_{2}\}$, where $\mathfrak{t}_{1}=\mathfrak{a}_{1}-\mathfrak{a}_{3}$ and $\mathfrak{t}_{2}=\mathfrak{a}_{2}-\mathfrak{a}_{3}$. Then the subcovering ${\mathcal{G}}_{\mathfrak{t}}={\mathcal{G}}/\Gamma_{\mathfrak{t}}$ of the diamond lattice ${\mathcal{G}}$ is the periodic graph shown in Fig. 3c.
@@ 1.3. Spectra of Schrödinger operators on subcovering graphs
Denote by $H_{\mathfrak{t}}=\Delta_{\mathfrak{t}}+Q$ the Schrödinger operator on the subcovering ${\mathcal{G}}_{\mathfrak{t}}={\mathcal{G}}/\Gamma_{\mathfrak{t}}$ with the potential $Q$ induced by the $\Gamma$-periodic potential $Q$ on the original periodic graph ${\mathcal{G}}$ (we use the same notation $Q$ for the induced potential). One can think of $H_{\mathfrak{t}}$ as of the Schrödinger operator $H=\Delta+Q$ on ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ acting on functions $f:{\mathcal{V}}\to{\mathbb{C}}$ that are periodic with respect to the sublattice $\Gamma_{\mathfrak{t}}\subset\Gamma$.
@@
$f(v+\mathfrak{t}_{s})=e^{i\langle t_{s},k\rangle}f(v)=f(v),\qquad\forall\,v\in{\mathcal{V}},\qquad\forall\,s\in{\mathbb{N}}_{d_{o}},$ (1.7)
where $\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}$, $d_{o}<d$, are the chiral vectors of the subcovering ${\mathcal{G}}_{\mathfrak{t}}$,
$\mathfrak{t}_{s}=t_{s1}\mathfrak{a}_{1}+\ldots+t_{sd}\mathfrak{a}_{d},\qquad t_{s}:=(t_{s1},\ldots,t_{sd})\in{\mathbb{Z}}^{d}.$ (1.8)
Let ${\mathcal{T}}$ be the $d_{o}\times d$ matrix given by
${\mathcal{T}}=\left(\begin{array}[]{ccc}t_{11}&\ldots&t_{1d}\\ \ldots&\ldots&\ldots\\ t_{d_{o}1}&\ldots&t_{d_{o}d}\\ \end{array}\right).$ (1.9)
The coordinate vectors $t_{s}\in{\mathbb{Z}}^{d}$ of the chiral vectors $\mathfrak{t}_{s}\in\Gamma$ with respect to the basis $\mathfrak{a}_{1},\ldots,\mathfrak{a}_{d}$ of the lattice $\Gamma$ will be called the chiral indices of the subcovering ${\mathcal{G}}_{\mathfrak{t}}$, and the matrix ${\mathcal{T}}$ whose rows are the chiral indices $t_{s}$, $s\in{\mathbb{N}}_{d_{o}}$, will be called the chiral matrix of ${\mathcal{G}}_{\mathfrak{t}}$.
@@
$\bullet$ we show that the spectra of the Schrödinger operators $H$ on a periodic graph ${\mathcal{G}}$ and $H_{\mathfrak{t}}$ on any its subcovering ${\mathcal{G}}_{\mathfrak{t}}$ with a primitive set $\mathfrak{t}$ of chiral vectors (i.e., a set $\mathfrak{t}$ which can be completed to a basis of the lattice $\Gamma$) consist of the same number of spectral bands, and the corresponding band edges are asymptotically close to each other as the chiral vectors are long enough (Proposition 2.2);
@@
We illustrate the obtained results by some examples of periodic graphs and their subcoverings. In particular, we describe all subcoverings of the $d$-dimensional lattice, the hexagonal lattice and the diamond lattice which are isospectral to the original periodic graphs (Examples 2.6, 2.8 and 2.10, respectively).

### v3

@@ 1. Introduction
This paper was partly motivated by the nice article [KP07] describing simple connections between spectra of Schrödinger operators on graphene and nanotubes. Graphene is a single 2D layer of graphite forming a hexagonal lattice. The hexagonal lattice is invariant under translations by vectors $\mathfrak{a}_{1}$ and $\mathfrak{a}_{2}$, see Fig. 1a. A carbon nanotube is a hexagonal lattice ”rolled up” into a cylinder. In carbon nanotubes, the graphene sheet is ”rolled up” in such a way that a so-called chiral vector $\mathfrak{t}=t_{1}\mathfrak{a}_{1}+t_{2}\mathfrak{a}_{2}$, where $t:=(t_{1},t_{2})\in{\mathbb{Z}}^{2}$, becomes the circumference of the tube, see Fig. 1b,c. In [KP07], applying a simple restriction procedure to the dispersion relation for the Schrödinger operator on the graphene lattice, the authors described the spectra of the Schrödinger operators on any carbon nanotube: zigzag ($t=(n,0)$), armchair ($t=(n,n)$) or chiral. In particular, they showed that the spectra of the Schrödinger operators on a nanotube and graphene coincide as sets if and only if $t_{1}-t_{2}$ is divisible by 3, and these spectra are asymptotically close as the length of the chiral vector $\mathfrak{t}$ tends to infinity. The similar approach was used in [D15] to derive spectral properties of graphyne nanotubes from the dispersion relation of the graphyne lattice. In [KS21] the authors also treated the connection between spectra of the Laplacians on a ${\mathbb{Z}}^{2}$-periodic 3-regular graph and its subcoverings to construct graphs with maximal gap intervals.
@@
Let ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ be a connected infinite graph embedded into ${\mathbb{R}}^{d}$. Here ${\mathcal{V}}$ is the set of its vertices and ${\mathcal{E}}$ is the set of its (unoriented) edges. An edge between vertices $u,v\in{\mathcal{V}}$ will be denoted as the (unordered) pair $\{u,v\}$ and said to be incident to the vertices $u$ and $v$.
@@
On edges of ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ we define a positive $\Gamma$-periodic weight
$\omega:{\mathcal{E}}\to(0,+\infty),\qquad\omega({\bf e}+\mathfrak{a})=\omega({\bf e}),\qquad\forall\,({\bf e},\mathfrak{a})\in{\mathcal{E}}\times\Gamma,$
and consider the Schrödinger operator $H$ acting on $\ell^{2}({\mathcal{V}})$ and given by
$H=\Delta+Q.$ (1.1)
Here $\Delta$ is the discrete weighted Laplacian having the form
$(\Delta f)(v)=\sum_{{\bf e}=\{v,u\}\in{\mathcal{E}}}\omega({\bf e})\big{(}f(v)-f(u)\big{)},\qquad f\in\ell^{2}({\mathcal{V}}),\qquad v\in{\mathcal{V}},$ (1.2)
@@
The sum in (1.2) is taken over all edges ${\bf e}\in{\mathcal{E}}$ incident to the vertex $v$. It is known that the Schrödinger operator $H$ is a bounded self-adjoint operator on $\ell^{2}({\mathcal{V}})$ (see, e.g., [SS92]).
We note that the embedding of the periodic graph ${\mathcal{G}}$ into ${\mathbb{R}}^{d}$ plays no role in the analysis of $H$, since only the incidence structure of ${\mathcal{G}}$ matters. Thus, without loss of generality we may assume that the period lattice $\Gamma$ of ${\mathcal{G}}$ is just the integer lattice ${\mathbb{Z}}^{d}$.
@@
$f(v+\mathfrak{a})=e^{i\langle\mathfrak{a},k\rangle}f(v),\qquad\forall\,v\in{\mathcal{V}},\qquad\forall\,\mathfrak{a}\in{\mathbb{Z}}^{d},$ (1.3)
@@
which, in particular, yields
@@
Some of $\lambda_{j}(\cdot)$ may be constant, i.e., $\lambda_{j}(\cdot)=\Lambda=\mathop{\mathrm{const}}\nolimits$, on some subset of ${\mathcal{B}}$ of positive Lebesgue measure. In this case the Schrödinger operator $H$ on ${\mathcal{G}}$ has the eigenvalue $\Lambda$ of infinite multiplicity. We call $\Lambda$ a flat band. Thus, the spectrum of the Schrödinger operator $H$ on a periodic graph ${\mathcal{G}}$ has the form (for more details see, e.g., [HN09])
@@
i) The first spectral band $\sigma_{1}(H)=[\lambda_{1}^{-},\lambda_{1}^{+}]$ is never a flat band and $\lambda_{1}^{-}=\lambda_{1}(0)$, see Theorem 2.7 in [SY23] and Theorem 1 in [SS92].
@@ 1.2. Subcovering graphs
Let ${\mathcal{G}}$ be a ${\mathbb{Z}}^{d}$-periodic graph. (Recall that we identify the period lattice $\Gamma$ of ${\mathcal{G}}$ with the integer lattice ${\mathbb{Z}}^{d}$). We fix some basis $\mathfrak{a}_{1},\ldots,\mathfrak{a}_{d}$ of ${\mathbb{Z}}^{d}$ (the periods of the graph ${\mathcal{G}}$). Below the coordinates of all vectors of ${\mathbb{Z}}^{d}$ will be given with respect to this fixed basis $\mathfrak{a}_{1},\ldots,\mathfrak{a}_{d}$.
Let $\Gamma_{\mathfrak{t}}$ be a sublattice of ${\mathbb{Z}}^{d}$ with a basis $\mathfrak{t}:=\{\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}\}\subset{\mathbb{Z}}^{d}$, where $d_{o}<d$. We consider a subcovering graph ${\mathcal{G}}_{\mathfrak{t}}$ of a ${\mathbb{Z}}^{d}$-periodic graph ${\mathcal{G}}$ obtained as the quotient of ${\mathcal{G}}$ with respect to the sublattice $\Gamma_{\mathfrak{t}}$, i.e.,
@@
The hexagonal lattice ${\bf G}$ is a ${\mathbb{Z}}^{2}$-periodic graph, the vectors $\mathfrak{a}_{1},\mathfrak{a}_{2}$ are the fixed periods of ${\bf G}$, see Fig. 1a. For non-zero vectors $\mathfrak{t}\in{\mathbb{Z}}^{2}$ the subcoverings ${\bf G}_{\mathfrak{t}}={\bf G}/{\mathbb{Z}}\mathfrak{t}$ of the hexagonal lattice ${\bf G}$ are known as nanotubes. For example, the nanotube ${\bf G}_{\mathfrak{t}}$ with the chiral vector $\mathfrak{t}=(2,1)$ is shown in Fig. 1b. The so-called zigzag nanotube ${\bf G}_{\mathfrak{t}}$, $\mathfrak{t}=(2,0)$, is presented in Fig. 1c.
@@
Figure 2. a) The cubic lattice ${\mathbb{L}}^{3}$; the blue vectors $\mathfrak{a}_{1},\mathfrak{a}_{2},\mathfrak{a}_{3}$ are the periods of ${\mathbb{L}}^{3}$; the fundamental cell $\Omega$ is shaded. The chiral vectors $\mathfrak{t}_{1}=\mathfrak{a}_{1}+\mathfrak{a}_{2}-\mathfrak{a}_{3}$ and $\mathfrak{t}_{2}=2\mathfrak{a}_{1}+\mathfrak{a}_{2}$ are marked in red color. b) The subcovering ${\mathbb{L}}^{3}_{\mathfrak{t}_{1}}$ of ${\mathbb{L}}^{3}$ with the chiral vector $\mathfrak{t}_{1}$ (the triangular lattice). c) The subcovering ${\mathbb{L}}^{3}_{\mathfrak{t}}$ of ${\mathbb{L}}^{3}$ with $\mathfrak{t}=\{\mathfrak{t}_{1},\mathfrak{t}_{2}\}$. d) The fundamental graph ${\mathbb{L}}^{3}_{*}$ of the cubic lattice ${\mathbb{L}}^{3}$ and any its subcovering ${\mathbb{L}}^{3}_{\mathfrak{t}}$ with a primitive set $\mathfrak{t}$ of chiral vectors.
@@
The cubic lattice ${\mathbb{L}}^{3}$ is a ${\mathbb{Z}}^{3}$-periodic graph, the vectors $\mathfrak{a}_{1},\mathfrak{a}_{2},\mathfrak{a}_{3}$ are the periods of ${\mathbb{L}}^{3}$, see Fig. 2a. The subcovering ${\mathbb{L}}^{3}_{\mathfrak{t}_{1}}={\mathbb{L}}^{3}/{\mathbb{Z}}\mathfrak{t}_{1}$ of the cubic lattice ${\mathbb{L}}^{3}$ with the chiral vector $\mathfrak{t}_{1}=(1,1,-1)$ is the triangular lattice, Fig. 2b. Let $\Gamma_{\mathfrak{t}}$ be the sublattice of ${\mathbb{Z}}^{3}$ with the basis $\mathfrak{t}:=\{\mathfrak{t}_{1},\mathfrak{t}_{2}\}$, where $\mathfrak{t}_{1}=(1,1,-1)$ and $\mathfrak{t}_{2}=(2,1,0)$. Then the subcovering ${\mathbb{L}}^{3}_{\mathfrak{t}}={\mathbb{L}}^{3}/\Gamma_{\mathfrak{t}}$ of the cubic lattice ${\mathbb{L}}^{3}$ is the ${\mathbb{Z}}$-periodic graph with the period $\mathfrak{a}_{1}$ shown in Fig. 2c. Note that this ${\mathbb{Z}}$-periodic graph can also be considered as the subcovering of the triangular lattice (which is a ${\mathbb{Z}}^{2}$-periodic graph with the periods $\mathfrak{a}_{1},\mathfrak{a}_{2}$) with the chiral vector $\mathfrak{t}_{2}=(2,1)$.
@@
c)$v_{2}$$v_{1}$${\mathcal{G}}_{\mathfrak{t}}$$\scriptstyle\mathfrak{a}_{1}$
@@
The diamond lattice ${\mathcal{G}}$ shown in Fig. 3a is a ${\mathbb{Z}}^{3}$-periodic graph, the vectors $\mathfrak{a}_{1},\mathfrak{a}_{2},\mathfrak{a}_{3}$ are the periods of ${\mathcal{G}}$. The subcovering ${\mathcal{G}}_{\mathfrak{t}}={\mathcal{G}}/{\mathbb{Z}}\mathfrak{t}$ of the diamond lattice ${\mathcal{G}}$ with the chiral vector $\mathfrak{t}=(1,1,-1)$ is shown in Fig. 3b. Note that this subcovering of ${\mathcal{G}}$ is isomorphic to the square lattice. Let $\Gamma_{\mathfrak{t}}$ be the sublattice of ${\mathbb{Z}}^{3}$ with the basis $\mathfrak{t}:=\{\mathfrak{t}_{1},\mathfrak{t}_{2}\}$, where $\mathfrak{t}_{1}=(1,0,-1)$ and $\mathfrak{t}_{2}=(0,1,-1)$. Then the subcovering ${\mathcal{G}}_{\mathfrak{t}}={\mathcal{G}}/\Gamma_{\mathfrak{t}}$ of the diamond lattice ${\mathcal{G}}$ is the ${\mathbb{Z}}$-periodic graph with the period $\mathfrak{a}_{1}$ shown in Fig. 3c.
@@ 1.3. Spectra of Schrödinger operators on subcovering graphs
Denote by $H_{\mathfrak{t}}=\Delta_{\mathfrak{t}}+Q$ the Schrödinger operator on the subcovering ${\mathcal{G}}_{\mathfrak{t}}={\mathcal{G}}/\Gamma_{\mathfrak{t}}$ with the potential $Q$ and the weight $\omega$ induced by the ${\mathbb{Z}}^{d}$-periodic potential $Q$ and weight $\omega$ on the original periodic graph ${\mathcal{G}}$ (we use the same notation $Q$ and $\omega$ for the induced potential and weight). One can think of $H_{\mathfrak{t}}$ as of the Schrödinger operator $H=\Delta+Q$ on ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ acting on functions $f:{\mathcal{V}}\to{\mathbb{C}}$ that are periodic with respect to the sublattice $\Gamma_{\mathfrak{t}}\subset{\mathbb{Z}}^{d}$.
@@
$f(v+\mathfrak{t}_{s})=e^{i\langle\mathfrak{t}_{s},k\rangle}f(v)=f(v),\qquad\forall\,v\in{\mathcal{V}},\qquad\forall\,s\in{\mathbb{N}}_{d_{o}},$ (1.7)
where $\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}\in{\mathbb{Z}}^{d}$, $d_{o}<d$, are the chiral vectors of the subcovering ${\mathcal{G}}_{\mathfrak{t}}$.
Let ${\mathcal{T}}$ be the $d_{o}\times d$ integer matrix whose rows are the chiral vectors $\mathfrak{t}_{s}$, $s\in{\mathbb{N}}_{d_{o}}$:
${\mathcal{T}}=\left(\begin{array}[]{ccc}t_{11}&\ldots&t_{1d}\\ \ldots&\ldots&\ldots\\ t_{d_{o}1}&\ldots&t_{d_{o}d}\\ \end{array}\right),\qquad\textrm{where}\qquad\mathfrak{t}_{s}=(t_{s1},\ldots,t_{sd})\in{\mathbb{Z}}^{d}.$ (1.8)
This matrix ${\mathcal{T}}$ will be called the chiral matrix of the subcovering ${\mathcal{G}}_{\mathfrak{t}}$.
@@
$\bullet$ we show that the spectra of the Schrödinger operators $H$ on a periodic graph ${\mathcal{G}}$ and $H_{\mathfrak{t}}$ on any its subcovering ${\mathcal{G}}_{\mathfrak{t}}$ with a primitive set $\mathfrak{t}$ of chiral vectors (i.e., a set $\mathfrak{t}$ which can be completed to a basis of the lattice ${\mathbb{Z}}^{d}$) consist of the same number of spectral bands, and the corresponding band edges are asymptotically close to each other as the chiral vectors are long enough (Proposition 2.2);
@@
We illustrate the obtained results by some examples of periodic graphs and their subcoverings. In particular, we describe all subcoverings of the $d$-dimensional lattice, the hexagonal lattice and the diamond lattice (with the unit edge weight) which are isospectral to the original periodic graphs (Examples 2.6, 2.8 and 2.10, respectively).